OpenStudy (anonymous):

How would you find the exact value of cos⁻1(0) without a calculator?

OpenStudy (amistre64):

let y = cos-1(0) and take the cos of each side

OpenStudy (anonymous):

But, without a calculator? How would you take the cos of each side then?

OpenStudy (amistre64):

then since trig is 98% memorization, youll have to remember what cos(y)=0 relates to

OpenStudy (amistre64):

y = cos-1(0), just cos each side to undo the cos^-1 cos(y) = 0

OpenStudy (anonymous):

So would the cos undo the arccos? Then leave cos(y)=0?

OpenStudy (anonymous):

And then y=pi/2 because cos pi/2=0? @amistre64